### v1

# Mitigating distribution shift in machine learning-augmented hybrid simulation

Abstract
We study the problem of distribution shift generally arising in machine-learning augmented hybrid simulation, where parts of simulation algorithms are replaced by data-driven surrogates. We first establish a mathematical framework to understand the structure of machine-learning augmented hybrid simulation problems, and the cause and effect of the associated distribution shift. We show correlations between distribution shift and simulation error both numerically and theoretically. Then, we propose a simple methodology based on tangent-space regularized estimator to control the distribution shift, thereby improving the long-term accuracy of the simulation results. In the linear dynamics case, we provide a thorough theoretical analysis to quantify the effectiveness of the proposed method. Moreover, we conduct several numerical experiments, including simulating a partially known reaction-diffusion equation and solving Navier-Stokes equations using the projection method with a data-driven pressure solver. In all cases, we observe marked improvements in simulation accuracy under the proposed method, especially for systems with high degrees of distribution shift, such as those with relatively strong non-linear reaction mechanisms, or flows at large Reynolds numbers.
keywords:
Machine learning, Distribution Shift, Regularization, Error Analysis, Fluid dynamics
{MSCcodes}
68T99, 65M15, 37M05
1 Introduction
Many scientific computational applications, such as computational fluid dynamics (CFD) and molecular dynamics (MD) can be viewed as dynamical system modeling and simulation problems, which is tackled by rigorous numerical tools with theoretical guarantee [45, 26]. However, in many cases a part of the simulation workflow, such as the Reynolds stresses in Reynolds-averaged Navier-Stokes equation (RANS) [1] and the exchange-correlation energies in density function theory used to compute force-fields that drives MD simulations [27], depends on models that are either expensive to compute, or even unknown in practice. One thus often resort to a hybrid simulation method, where the known, resolved components of the dynamics are computed exactly, while the unresolved components are replaced by approximate, but computationally tractable models. For example, solving Navier-Stokes equations using projection method involves two steps. In the first step, all the terms except for the gradient of the pressure is used to evolve the velocity. This step is computationally cheap and thus understood as the resolved part. Next, the pressure is solved from a Poisson equation and then used to correct the velocity. Most of the computational cost is contained in solving this Poisson equation and we thereby viewed this as the unresolved part. We call such scientific computing problems with both resolved and unresolved parts “hybrid simulation problems”. Similar problems are surveyed in [51] under the name of “Hybrid physics-DL models”.
As machine learning becomes increasingly powerful in areas like computer vision and natural language processing, practitioners begin to use data-driven modules to model the unresolved part to carry out the simulation. For example, in [49] the author replaced the numerical Poisson solver of the unresolved part by a convolutional neural network trained using a novel unsupervised learning framework. Then, this data-driven model is coupled with resolved model and provide fast and realistic simulation results in 2D and 3D. Similar ideas are used else where, e.g. RANS [28]. We hereafter refer to this simulation procedure by machine-learning augmented hybrid simulation (MLHS). This structure represents a great part of the scientific machine learing research [54, 46, 25] and will be the focus of this paper. In MLHS, a common problem appears: while the data-drive model performs well on the training data, the performance quickly deteriorates when iteratively applying it in simulation, driving the dynamics to some regime which is not observed from the training source. Empirical evidences have been observed in various applications, such as CFD [46, 35, 9, 54], molecular dynamics [56, 55, 25], and iterative numerical solver [2] but these problems are not well-studied algorithmically and theoretically in the literature.
This issue has strong connections with the so-called distribution shift (DS) in computer science applications, especially in reinforcement learning and computer vision. Researchers use DS to refer to problems where training and testing distribution differ significantly. An example is an image classifier trained with images taken in the daytime that is tested under night conditions [23]. Therefore, accuracy on the training regime itself cannot guarantee the performance during inference. To resolve this, researchers have systematically developed several methods, e.g. domain adaptation [11], ensemble learning [8], and meta-learning [50]. However, there are key differences between the distribution shift phenomena in MLHS and that in traditional applications between the DS phenomena in MLHS and computer vision. While distribution shift in computer vision lacks a theoretical model to describe [23], that in MLHS comes from dynamical systems for which we have abundant knowledge on the resolved parts of the models. For example, let us consider using projection method [10] to solve the incompressible Navier-Stokes equation. Suppose we replace the Poisson pressure solver by a data-driven model based on neural network. Then, the dynamics of this data-driven hybrid numerical solver will largely depend on the properties of the resolved numerical part of the projection method, which is well-studied [10]. By analyzing the stability properties of this dynamics, one can quantify which family of distribution shifts may arise and resolve them according to the information of the dynamics. Therefore, one can pose the following question: Can we use the information of resolved parts to design robust learning algorithm for unresolved parts to improve hybrid simulation?
In this paper, we first develop a mathematical framework to understand the origin of distribution shift in MLHS and how it may lead to simulation performance deterioration. We emphasize the difference between this instability issue in data-driven scientific computing and distribution shift in the machine learning literature, such as reinforcement learning and computer vision. Then, we propose an algorithm to improve the simulation accuracy by mitigating distribution shift. We assume by manifold hypothesis [34] that the correct trajectories lie on a low-dimensional manifold of the high dimensional space, e.g. the fluid configuration of Navier-Stokes equation lies on the solution manifold of the high-dimensional grid space. The key idea is to combine the physical information of the resolved part of the model, i.e. Navier-Stokes equation and this manifold structure learned from the data to form a regularization term for the data-driven model. Intuitively speaking, this regularizer stablizes the dynamics by preventing it from moving further away the data manifold. Such movements may result in configurations that are either non-physical, or corresponds to different initial/boundary conditions. Therefore, preventing such movements reduces the severity of distribution shift, and the data-driven model will stay relatively accurate during the simulation, which promotes high simulation fidelity in a long time interval. In implementation, we first use an autoencoder (AE) to parameterize the underlying data manifold. After this preprocessing, the AE is combined with the resolved dynamics and plugged into the loss function of the model as a regularization, which prevents the simulation from moving to unseen regimes. One then back-propagates this modified loss function to optimize the data-driven model. The algorithm is tested on several representative numerical examples to demonstrate its effectiveness. We show that the proposed approach can improve simulation accuracy under different extents of distribution shift. Indeed, the improvements become more significant in scenarios where the fidelity of the simulation is highly sensitive to errors introduced in the data-driven surrogates, such as fluid simulations at high Reynolds numbers and reaction-diffusion equations with relatively strong non-linear reaction. Specifically, in high Reynolds number fluid simulations, naive data-driven surrogate models may quickly cause error blow-ups, but our method can maintain simulation accuracy over large time intervals.
The paper is structured as follows. In Section 2, we establish a precise framework to identify the origins of distribution shift in MLHS. In Section 3, we introduce our regularized learning algorithm motivated from the analysis of the previously identified form of distribution shift and theoretically understand its performance in the linear setting. We also discuss connections with the literature on control and system identification. In Section 4, we validate our algorithm on several practical numerical cases, including simulating a reaction-diffusion equation and the incompressible Navier-Stokes equation.
2 Mathematical formulation of MLHS and the problem of distribution shift
In this section, we provide the mathematical formulation of MLHS and then identify the problem of distribution shift. In the first subsection we give a general treatment, and we provide concrete examples in Section 2.2.
2.1 The resolved and unresolved components in hybrid simulations
Throughout this paper, we consider the dynamics
$\partial_{t}\mathbf{u}$ $=\mathcal{L}(\mathbf{u},\mathbf{y},t),\quad\mathbf{u}\in\mathbb{R}^{m},\mathbf{y}\in\mathbb{R}^{n},\mathcal{L}:\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{m},$ (1)
$\mathbf{y}$ $=\phi(\mathbf{u},t),\quad\phi:\mathbb{R}^{m}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{n},$
where $\mathbf{u}$ is the resolved state variable, e.g. fluid velocity field or chemical concentration field. The vector $\mathbf{y}$ is the unresolved variable, which drives the dynamics for $\mathbf{u}$ but is either expensive to compute or cannot be directly observed. Throughout the paper, we will omit writing the explicit time-dependence by using $\mathbf{u}$ ($\mathbf{y}$) to denote $\mathbf{u}(t)$ ($\mathbf{y}(t)$) whenever there is no ambiguity. We also restrict the discussion to the finite dimension case, but most of the analysis here will be applicable to the more general case of $\mathbf{u},\mathbf{y}$ belonging to an infinite dimensional Hilbert space. We adopt the following assumptions associated with Eq. 1
  • 1. Resolved component: $\mathcal{L}$ is known, possibly non-linear.
  • 2. Unresolved component: $\phi$ is either unknown, partially-known or expensive to evaluate.
Given such a system, the goal is to first obtain information on the unresolved model and then integrate it with resolved part to simulate the whole dynamics. This hybrid structure, which we call hybrid simulation problems is general enough to include most of the settings in hybrid simulation, e.g. when simulating the trajectories of molecular dynamics which satisfy some stochastic differential equations or the time-evolution of fluid velocity fields following the Navier-Stokes equation. This paper focuses on MLHS, a particular variation of such hybrid simulation problem where the unresolved component is tackled by a data-driven method. This point will be made more precise in Section 2.3. To simplify the analysis, we adopt the forward Euler time discretization for simulation
$\widehat{\mathbf{u}}_{k+1}$ $=\widehat{\mathbf{u}}_{k}+\Delta t\mathcal{L}(\widehat{\mathbf{u}}_{k},\widehat{\mathbf{y}}_{k}),\qquad\widehat{\mathbf{y}}_{k}=\phi(\widehat{\mathbf{u}}_{k}),$ (2)
while other consistent discretizations may be analyzed in the same spirit. Here we drop the dependence of $\mathcal{L},\phi$ on $t$ to consider autonomous system while our later discussion and algorithm are also applicable to non-autonomous system.
2.2 Examples of hybrid simulation problems
Before moving on to distribution shift, we first provide some examples in MLHS of our interest, which are also closely related to our numerical experiments. We will present these examples according to the structure Eq. 1.
The first example is solving the 2D incompressible Navier-Stokes equation using the projection method [17, 10], a variant of which is as follows:
$\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}-\nu\Delta\mathbf{u}=\nabla p,\quad\nabla\cdot\mathbf{u}=0,\quad T\in[0,1],$ (3)
where $\mathbf{u}=(u(x,y,t),v(x,y,t))^{T}\in\mathbb{R}^{2}$ is velocity and $p$ pressure. Fix a regular grid size for discretization. The projection method can be written as follows
$\mathbf{u}_{k+1}$ $=\mathbf{u}_{k}+\Delta t\mathcal{L}(\mathbf{u}_{k},p_{k})$ (4)
$=\mathbf{u}_{k}+\Delta t(\nu\Delta\mathbf{u}_{k}-(\mathbf{u}_{k}\cdot\nabla)\mathbf{u}_{k}-\nabla p_{k}),$
$p_{k}$ $=\phi(\mathbf{u}_{k})=\Delta^{-1}(\nabla\cdot\left(\nu\Delta\mathbf{u}_{k}-(\mathbf{u}_{k}\cdot\nabla)\mathbf{u}_{k}\right)),$
We write $\Delta^{-1}$ as the inverse operator for Poisson equation formally. The resolved part is performed by first stepping the convection and diffusion term then using pressure to correct the step, and the unresolved part is a Poisson equation which related the unsolved state $p_{k}$ with the velocity $\mathbf{u}_{k}$. Here, $p_{k}$ represents the unresolved variable $\mathbf{y}_{k}$ in our formulation Eq. 1.
In this method, the most expensive step is the pressure computation, which requires repeated solutions of similar large-scale linear equations. State of the art solvers such as multigrid [4] and conjugate gradient [43] become prohibitively expensive when the problem size is large. Recently, a promising direction is to replace the Poisson solver by data-driven surrogate models. Ref. [49] replaces the pressure calculation step by a convolutional neural network trained with unsupervised learning by requiring the updated velocity to have zero divergence. Similarly, machine learning models such as tensor-based neural networks are used in [28] to replace classical turbulence modeling for unclosed Reynolds stresses tensor. This model is then plugged into the Reynolds-averaged Navier-Stokes simulation to predict the flow separation. More recently, [37] combines the idea of reduced-order modeling with graph convolutional neural network to encode the reduced manifold and enable fast evaluations of parametrized PDEs.
The second example is based on simulation on two grids of different sizes. Let us continue with the projection method setting Eq. 4. Denote the time-evolution operator on grid size $n$ as $f_{n}:\mathbf{u}_{k}^{n}\rightarrow\mathbf{u}_{k+1}^{n}$, which becomes very expensive to compute when $n$ is large. Therefore, one may wish to replace the fine-grid solver with a coarse-grid one and add some correction. This again reduces the computational burden for high-fidelity simulations. This structure shows up in various scientific computing situations, such as multigrid solvers [4], Reynolds-averaged Navier-Stokes (RANS) [38, 1], and large eddy simulation (LES) [59]. Suppose we have a fine grid with size $2n\times 2n$ and a coarse grid with size $n\times n$, we use superscript $n,2n$ to denote the field defined on these two grids, i.e. $\mathbf{u}_{k}^{n},\mathbf{u}_{k}^{2n}$. Moreover, we fix an interpolation and a restriction operator between fields on these two grids:
$I_{n}^{2n}:\mathbb{R}^{n\times n}\rightarrow\mathbb{R}^{2n\times 2n},\quad R_{2n}^{n}:\mathbb{R}^{2n\times 2n}\rightarrow\mathbb{R}^{n\times n}.$ (5)
Now, we can state the second hybrid simulation problem as follows:
[…]
In detail, the resolved component contains a solver of the evolution equation over the coarse grid which marches forward by one time step, with field given by a restriction of the field on the fine grid. Then, the next step field configuration is interpolated to the fine grid, and the unresolved component serves to correct the deviation of the field variables between simulating on grids of different sizes. The unresolved variables $\mathbf{y}_{k}^{2n}$ can be calculated if we have accurate simulation results $\mathbf{u}_{k}^{2n}$ on the fine grid by
$\mathbf{y}_{k}^{2n}=\mathbf{u}_{k+1}^{2n}-I_{n}^{2n}\circ f_{n}(R_{2n}^{n}(\mathbf{u}_{k}^{2n}))).$ (7)
In the literature, [21] treats the smoothing algorithm in multigrid solver as unresolved components and other procedures such as restriction and prolongation as resolved components. This fits into our framework of MLHS. They introduce a supervised loss function based on multigrid convergence theory to learn the optimal smoother and improve the convergence rate over anisotropic rotated Laplacian problems and variable coefficient diffusion problems. In [32], the authors applies neural network as non-linear regression to fit between some key fluid features and sub-grid stresses. Going beyond this, [35] adds neural emulation to offline learning to prevent the trajectories from deviating away the ground truth. This data-driven model is plugged into hybrid simulation to test its ability of preserving coherent structure and scaling laws.
2.3 Learning the unresolved model and distribution shift
We assume that the resolved part $\mathcal{L}$ is known, or we at least have access to a gray box that can perform its evaluation and compute its gradients. While the unresolved part is unknown, we can access data tuples $\left\{(\mathbf{u}_{1},\mathbf{y}_{1},t_{1}),(\mathbf{u}_{2},\mathbf{y}_{2},t_{2}),\cdots,(\mathbf{u}_{N},\mathbf{y}_{N},t_{N})\right\}$ which are obtained either by physical experiments or accurate but expensive numerical simulations. The goal is hence to construct an approximate mapping $(\mathbf{u}_{k},t_{k})\mapsto\mathbf{y}_{k}$ from this dataset. Viewing this as a supervised learning task, there are many existing methods based on empirical risk minimization [19, 20, 24]. One of the simplest method is based on minimizing the $L_{2}$ loss function. Given a parameterized model $\phi_{\theta}$, either a classical one with unknown parameters or a surrogate one such as a neural network, we learn from data the optimal value of the parameter $\theta$, i.e. $\widehat{\theta}=\arg\min_{\theta}\mathbb{E}_{(\mathbf{u},\mathbf{y})}\left\|\mathbf{y}-\phi_{\theta}(\mathbf{u},t)\right\|^{2}$. However, we emphasize that the data assumption in learning of dynamical processes differs significantly with the traditional i.i.d. assumption in statistical learning [20]. As the data is composed by several simulation trajectories, those belonging to the same trajectory will have large correlation. Therefore, it is reasonable to expect that simple empirical risk minimization (which many existing MLHS works employ [46, 42]) may not be the optimal method. We will demonstrate this by using the least squares estimator as a baseline comparison to our proposed approach Section 4.
After determining the parameters of the unresolved model, one can perform hybrid simulation to obtain new trajectories, i.e.
$\widehat{\mathbf{u}}_{k+1}=\widehat{\mathbf{u}}_{k}+\Delta t\mathcal{L}(\widehat{\mathbf{u}}_{k},\widehat{\mathbf{y}}_{k},t_{k}),\quad\widehat{\mathbf{y}}_{k}=\phi_{\widehat{\theta}}(\widehat{\mathbf{u}}_{k},t_{k}).$ (8)
Notice we add hat superscript to all the quantities related to this simulated dynamics to distinguish them from the ground truth $(\mathbf{u}_{k},\mathbf{y}_{k})$. We measure the performance of the simulator by the error along the whole trajectories, i.e.
$\min\sum_{k=1}^{n}\left\|\mathbf{u}_{k}-\widehat{\mathbf{u}}_{k}\right\|^{2},$ (9)
where $\mathbf{u}_{k}$s are the true trajectories and $\widehat{\mathbf{u}}_{k}$s are the simulated one with the same initial condition.
We begin by illustrating the issue of distribution shift by analyzing of the trajectory error. For simplicity of presentation, we assume that the model hypothesis space $\{\phi_{\theta}:\theta\in\Theta\}$ is bias-free, meaning that there exists $\theta^{*}\in\Theta$ such that $\phi=\phi_{\theta^{*}}$. In the general case where the hypothesis space is universal but not closed, this equality would be replaced by an approximate one, but the argument follows analogously. Comparing Eq. 1 and Eq. 8, we have
$\ \widehat{\mathbf{u}}_{k+1}-\mathbf{u}_{k+1}$ (10)
$=$ $\ \widehat{\mathbf{u}}_{k}-\mathbf{u}_{k}+\Delta t\left(\mathcal{L}(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))-\mathcal{L}(\mathbf{u}_{k},\phi_{\theta_{*}}(\mathbf{u}_{k}))\right)$
$+\Delta t\left(\mathcal{L}(\widehat{\mathbf{u}}_{k},\phi_{\widehat{\theta}}(\widehat{\mathbf{u}}_{k}))-\mathcal{L}(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))\right).$
The error introduced by moving one step forward is thus decomposed into two parts: the error associated with the estimator $\widehat{\theta}:\mathcal{L}(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))-\mathcal{L}(\mathbf{u}_{k},\phi_{\theta_{*}}(\mathbf{u}_{k}))$, and the other error $\mathcal{L}(\widehat{\mathbf{u}}_{k},\phi_{\widehat{\theta}}(\widehat{\mathbf{u}}_{k}))-\mathcal{L}(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))$. The former resembles the error appearing in classical statistical inference problem under the i.i.d. setting since $\mathbf{u}_{k}$ follows an identical distribution of the training dataset. However, the latter is different in the sense that $\widehat{\mathbf{u}}_{k}$ and $\mathbf{u}_{k}$ may come from distributions of their respective driving dynamics, which are different since $\theta_{*}\neq\widehat{\theta}$. This may be due to noisy labels, few data points, or having an under-determined system. This difference between MLHS and statistical learning is also emphasized in [51]. This latter error is rather akin to the issue of stability in numerical solution for partial differential equations: although the error in each step is relatively small, it can accumulate exponentially as time step iterates, driving $\widehat{\mathbf{u}}_{k}$ to completely different regime where we have not observed in the data $\mathbf{u}_{k},k=1,\cdots,N$. As a result, the data-driven model $\widehat{\phi}$ can no longer be trusted to be accurate, since it is trained on data from a different distribution. This may then lead to a vicious cycle, where further deterioration of trajectory error occurs, leading to increased discrepancy between the distributions. Consequently, one can understand the distribution shift as a large discrepancy between the true distribution of $\mathbf{u}_{k}$ and simulated distribution $\widehat{\mathbf{u}}_{k}$. This problems occur commonly in MLHS. For example, [28, 29, 53, 39] all observe unphysical solutions when substituting Reynolds stresses models calculated from direct numerical simulation database into RANS. Ref. [53] attributes this to the ill-conditioning of the RANS problems and propose a method to treat Reynolds stresses implicitly, while [30] uses coupling modes perspective to numerically study the stability and convergence of RANS and [18] derives the propagation error of RANS simulation error mathematically. However, a general theoretical framework to understand and combat distribution shift in MLHS is lacking. The current paper aims to make progress in this direction.
Let us illustrate the problem of distribution shift using numerical examples. We solve the 2D FitzHugh-Nagumo reaction-diffusion equation over a periodic domain $[0,6.4]\times[0,6.4]$, whose dynamics is given by
$\frac{\partial\mathbf{u}}{\partial t}$ $=D\Delta\mathbf{u}+\phi(\mathbf{u}),\quad T\in[0,20],$ (11)
$\phi(\mathbf{u})$ $=\phi(u,v)=\begin{pmatrix}u-u^{3}-v+\alpha\\ \beta(u-v)\end{pmatrix}.$
where $\mathbf{u}=(u(x,y,t),v(x,y,t))^{T}\in\mathbb{R}^{2}$ are two interactive components, $D$ is the diffusion matrix, and $\mathbf{R}(\mathbf{u})$ is source term for the reaction. We fix the parameters to $\alpha=0.01,\beta=1.0,D=\begin{pmatrix}0.05&0\\ 0&0.1\end{pmatrix}$, where the system is known to form Turing patterns [33]. Assuming that we do not know the exact form of the reaction term $\phi(\mathbf{u})$, the resolved part is the diffusion term and the unresolved part is the non-linear reaction term. The ground truth data are calculated using the semi-implicit Crank-Nicolson scheme with full explicit discretization for the non-linear term. As we do not know the reaction term is pointwise in space, we use a convolutional neural network-based model to learn it based on ordinary least squares (OLS). Specifically, we solve the least squares problem $\min_{\theta}\mathbb{E}_{\mathbf{u}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|^{2}$ using Adam [22]. Moreover, we conduct an ablation study where the unresolved model consists of the exact form of the non-linear term as in Eq. 11, plus a random Gaussian noise whose variance is set equal to the optimization error of the OLS estimator, which is $10^{-4}$ in the following experiments. By comparing this and the simulation based on the OLS model, we can conclude whether or not the trajectory error is caused mainly by the estimation error of the unresolved model, or by the shift in distribution of the input data fed to the unresolved model. More details on the data generation and optimization procedure is described in the supplementary materials LABEL:numeric-setup. We simulate this PDE with a test initial condition sampled from the same distribution as training trajectories and compare the results with the ground truth. We present the comparison in Fig. 1.
Figure 1: The first three rows display the snapshots of ground truth, simulated fields, and ablation study simulation at time steps: 0, 100, 200, 300, 500. (d) plots the relative error and distribution shift between two flow configurations at corresponding time. The error is defined as $\frac{\left\|\widehat{\mathbf{u}}_{t}-\mathbf{u}_{t}\right\|_{2}}{\left\|\mathbf{u}_{t}\right\|_{2}}$ at each time step $t$ instead of the whole trajectory as in Eq. 9. The plotted distribution shift roughly measures the average discrepancy between the distributions of $\mathbf{u}_{k}$ and $\widehat{\mathbf{u}}_{k}$ at time step $k$, and is calculated using an autoencoder explained in Section 3.2. Although the OLS estimator and ablation-study estimator share the same error magnitude for unresolved model, their long time trajectory performance form sharp contrast, indicating the optimization error of the unresolved model could not explain the failure in long time trajectory prediction.
The first three rows display the snapshots of ground truth, simulated fields, and ablation study simulation at time steps: 0, 100, 200, 300, 500. (d) plots the relative error and distribution shift between two flow configurations at corresponding time. The error is defined as $\frac{\left\|\widehat{\mathbf{u}}_{t}-\mathbf{u}_{t}\right\|_{2}}{\left\|\mathbf{u}_{t}\right\|_{2}}$ at each time step $t$ instead of the whole trajectory as in Eq. 9. The plotted distribution shift roughly measures the average discrepancy between the distributions of $\mathbf{u}_{k}$ and $\widehat{\mathbf{u}}_{k}$ at time step $k$, and is calculated using an autoencoder explained in Section 3.2. As can be observed, it does not take a very long time for the distribution shift issue to be severe enough using ordinary least squares estimator. Comparing the OLS error and distribution shift with the those of the ablation study, we conclude that the estimation error is not the main cause of the trajectory error, since the ablation-study trajectory has the same magnitude of estimation error, but has much smaller trajectory error and distribution shift. Moreover, in both cases the relative error of the trajectory has the same trend as the distribution shift, indicating correlation.
Indeed, this phenomenon is ubiquitous in modern MLHS [51, 56, 46]. As a second example, we solve the Navier-Stokes equation Eq. 3 using projection method Eq. 4 where pressure solver is replaced by a learned convolutional neural network predictor. The predictor is optimized over data pair of fluid velocity and pressure obtained from high-fidelity simulation to attain an error of $10^{-5}$. Due to the complexity of projection solver and grid issue, we do not conduct ablation study in this case.
Figure 2: Configurations of the Navier-Stokes equation simulation with Reynolds number $200$ at time step: 0, 60, 120, 180, 240: (a) ground truth (b) simulated fluid field using OLS estimator. We only show velocity field over part of the domain. The dynamics blow up in the last configuration, which does not show up in the case of reaction-diffusion equation. This can be explained by the fact that the Navier-Stokes equation is more unstable than the reaction-diffusion equation. In (c), we illustrate the error along the trajectories, the distribution shift.
The distribution shift in Fig. 2 turns out to be more severe when we simulate the incompressible Navier-Stokes equation. After $160$ time steps, the fluid configuration becomes rough and un-physical. Here, we again observe that the DS and error increase together, suggesting correlation. In the next section, we will make this connection precise, and furthermore develop principled methods to control the distribution shift, thereby improving the prediction fidelity.
3 Theoretical analysis and algorithm
In this section, we first put the distribution shift issue into a theoretical framework. Then, we introduce and analyze our algorithm in this section. We will provide rigorous study of the linear case and use this to motivate the algorithm.
3.1 Distribution shift in linear dynamics and motivation of tangent-space regularized estimator
As there exists few tools to tackle the general non-linear dynamics Eq. 1, we switch to the simpler case of linear dynamics, which also capture some key features in general situations. We consider the following hybrid simulation problem
$\partial_{t}\mathbf{u}$ $=A\mathbf{u}+B\mathbf{y},\quad\mathbf{u}\in\mathbb{R}^{m},\mathbf{y}\in\mathbb{R}^{n},A\in\mathbb{R}^{m\times m},B\in\mathbb{R}^{m\times n}$ (12)
$\mathbf{y}$ $=C^{*}\mathbf{u},\quad C^{*}\in\mathbb{R}^{n\times m}.$
Such structures appear in many scenarios, e.g. finite difference solution of discretized linear PDEs and linear control problems. Alternatively, one can think of this linear system as the linearization of the non-linear dynamics Eq. 1 about some steady-state.
We assume that the state variables $\mathbf{u}_{i}$s in training data come from several trajectories and $\mathbf{y}_{i}$s may be subject to some random measurement error, i.e. $\mathbf{y}_{i}=C^{*}\mathbf{u}_{i}+\epsilon_{i}$, or in matrix form $Y=C^{*}\mathbf{U}+\epsilon$. The key point is that these trajectories may belong to a low-dimensional subspace in the high-dimensional state space. This is the case for many scientific computing problems, such as computational fluid dynamics [38], and is also known more generally as the manifold hypothesis [34]. We use $V\subset\mathbb{R}^{m}$ to denote the subspace that contains all the training data $\mathbf{u}_{i}$. In this setting, the appearance of this low-dimensional structure may be caused by two different situations. In the first situation, the initial value of $\mathbf{u}$ is supported on the subspace spanned by several eigenvectors of the evolution operator $e^{(A+BC^{*})}$, i.e.
$\mathbf{u}_{0}\in\operatorname{span}\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{l}\},\quad e^{(A+BC^{*})}\mathbf{v}_{1}=\lambda_{1}\mathbf{v}_{1},\cdots,e^{(A+BC^{*})}\mathbf{v}_{l}=\lambda_{l}\mathbf{v}_{l},$ (13)
Then, all the training data $\mathbf{u}_{i}$ belong to this subspace, i.e. $V=\operatorname{span}\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{l}\}$. The second situation is that the dynamics is degenerate in the sense that the evolution matrix $A$ is not of full rank. This will not appear in the differential formulation as infinitesimal transformations in the form $e^{(A+BC^{*})\Delta t}$ are always non-degenerate. However, for a discrete dynamics
$\mathbf{u}_{k+1}=A\mathbf{u}_{k}+B\mathbf{y}_{k},\quad\mathbf{y}_{k}=C^{*}\mathbf{u}_{k},$ (14)
if the matrix $A+BC^{*}$ is degenerate, then all the data $\mathbf{u}_{k}$ will belong to the range of this operator $A+BC^{*}$, which is a proper subspace. Most cases of scientific computing applications belong to the first class. Due to the degeneracy in the training data, the least squares solution is not unique. Hence, various empirical regularizations (e.g. $\ell^{2}$ regularization) can be introduced to obtain an estimator with desirable properties.
Now, suppose one obtains an estimator $\widehat{C}$, one can calculate several error metrics. The first is the test error
$l_{OLS}(\widehat{C})=\mathbb{E}\left\|(\widehat{C}-C^{*})\mathbf{u}\right\|^{2},$ (15)
where $\mathbf{u}$ is sampled from the low dimensional subspace $V$ and follows the same distribution as the training data. We refer to this as statistical estimation error. However, we are actually interested in the error of the simulated dynamics
$\partial_{t}\widehat{\mathbf{u}}=A\widehat{\mathbf{u}}+B\widehat{\mathbf{y}},\quad\widehat{\mathbf{y}}=\widehat{C}\widehat{\mathbf{u}},$ (16)
compared to the ground truth $\mathbf{u}$ with the same initial condition. The time evolution of their difference is given by
$\partial_{t}(\widehat{\mathbf{u}}-\mathbf{u})=$ $\ (A+B\widehat{C})\widehat{\mathbf{u}}-(A+BC^{*})\mathbf{u}$ (17)
$=$ $\ A(\widehat{\mathbf{u}}-\mathbf{u})+B(\widehat{C}\widehat{\mathbf{u}}-\widehat{C}P_{V}\widehat{\mathbf{u}}+\widehat{C}P_{V}\widehat{\mathbf{u}}-\widehat{C}\mathbf{u}+\widehat{C}\mathbf{u}-C^{*}\mathbf{u})$
$=$ $\ (A+B\widehat{C}P_{V})(\widehat{\mathbf{u}}-\mathbf{u})+B(\widehat{C}-C^{*})\mathbf{u}+B\widehat{C}(\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}}),$
where $P_{V}$ is the orthogonal projection onto the data subspace $V$ and we have $P_{V}\mathbf{u}=\mathbf{u}$ since $\mathbf{u}$ belongs to the data subspace. As derived in Eq. 17, the overall error of the trajectory during evolution can be decomposed into three parts. In the following we make a detailed analysis of each term.
The first term can be viewed as an amplitude factor of the error propagation and can be easily bounded by
$\left\|(A+B\widehat{C}P_{V})(\widehat{\mathbf{u}}-\mathbf{u})\right\|_{2}\leq\left\|A+B\widehat{C}P_{V}\right\|_{2}\left\|\widehat{\mathbf{u}}-\mathbf{u}\right\|_{2}.$ (18)
The operator $A+B\widehat{C}P_{V}$ can be viewed as a posterior estimation of the error propagation and a stability constant of the simulated dynamics.
The second term is simply the statistical estimation error of the estimator $\widehat{C}$ defined in Eq. 15. Given that the initial condition $\mathbf{u}_{0}$ of the training and test trajectories are sampled from the same distribution, this term is well-bounded in classical statistical learning theory [20], provided the number of data tuples is sufficiently large. In particular, this error has nothing to do with distribution shift.
Now we move to the last part $B\widehat{C}(\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}})$, which can be more easily understood from a geometric viewpoint. The last factor $(\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}})$ measures how far the simulated trajectory is away from the data subspace $V$. In other words, this term measures how far the simulated trajectory’s state distribution “shifts” from the true distribution. This is exactly the term we want to obtain to control the distribution shift in simulation. Meanwhile, this term forms sharp contrast with previous two terms in the sense that it is not automatically bounded in most algorithm for estimating $\widehat{C}$. We will illustrate this point further by deriving a error bound of the simulated dynamics, which shows that it depends sensitively on this distribution shift term. Thus, we can improve prediction accuracy if we combat such distribution shifts.
Now, taking the expectation of the inner product between $\widehat{\mathbf{u}}-\mathbf{u}$ and Eq. 17 and applying Gronwall’s lemma one obtains the following bound on the error propagation.
$\mathbb{E}\left\|\widehat{\mathbf{u}}(T)-\mathbf{u}(T)\right\|_{2}^{2}\leq$ (19)
$\frac{\left\|B\right\|_{2}^{2}\mathbb{E}\left\|(\widehat{C}-C^{*})\mathbf{u}\right\|^{2}+\left\|B\widehat{C}\right\|_{2}^{2}\sup_{t\leq T}\left\|\widehat{\mathbf{u}}(t)-P_{V}\widehat{\mathbf{u}}(t)\right\|_{2}^{2}}{\left\|A+B\widehat{C}P_{V}\right\|_{2}^{2}+3}(e^{(\left\|A+B\widehat{C}P_{V}\right\|_{2}^{2}+3)T}-1).$
An interpretation for this result is that the trajectory error can be divided into two parts: $\left\|B\right\|_{2}^{2}\mathbb{E}\left\|(\widehat{C}-C^{*})\mathbf{u}\right\|^{2}$ and $\left\|B\widehat{C}\right\|_{2}^{2}\sup_{t}\left\|\widehat{\mathbf{u}}(t)-P_{V}\widehat{\mathbf{u}}(t)\right\|_{2}^{2}$. The first error is the statistical estimation error of unresolved component, which can be reduced by better training algorithms, better models, or more data etc. This is the familiar term one faced in supervised learning under common i.i.d. setting. The second error relates to the distribution shift since it is measured by the difference between the sampled trajectories and underlying data manifold, i.e. $\left\|\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}}\right\|=\operatorname{dist}(\widehat{\mathbf{u}},V)$. Moreover, this shift is caused by the interaction between the estimator of unresolved model and the resolved model dynamics. Specifically, the estimator makes some error in prediction and this error is propagated via the resolved model dynamics and then fed into the next round of hybrid simulation. This distribution shift in MLHS is far different from the one in computer science such as covariate shift and label shift. In computer vision, the cause of distribution shift may be extrinsic [23], i.e. the pictures in training set are all taken at daytime while those for testing are all taken at night. Such shifts are hard to model, so their resolution tends to depend on data augmentation and related techniques, e.g. the Dagger algorithm [41]. Instead of changing the optimization problem, the data source is modified. This shares much similarity with adversarial training, which also add more data to the training set to make the prediction robust under adversarial attack, a typical distribution shift in the area of computer vision [15]. In contrast, in our setting the distribution shift is intrinsically driven by the hybrid simulation structure of which we have partial knowledge. Hence, we can quantify this distribution shift and design specific algorithm to mitigate them.
Returning to the linear problem, in order to guarantee that the error $\widehat{\mathbf{u}}-\mathbf{u}$ is bounded over the simulated trajectory, a natural choice is to use a regularization for $\left\|\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}}\right\|_{2}$. The most naive choice would be $\left\|\widehat{\mathbf{u}}-P_{V}\widehat{\mathbf{u}}\right\|_{2}$ itself, but a problem is that $\widehat{\mathbf{u}}$ is calculated via hybrid simulation until time $t$, which corresponds to a rather complicated computational graph. This makes subsequent gradient-based optimization computational expensive. Hence, we take the advantage of an one-step predictor, i.e. we set
$\widehat{\mathbf{u}}_{k+1}=\mathbf{u}_{k}+\Delta t(A\mathbf{u}_{k}+B\widehat{C}\mathbf{u}_{k}).$ (20)
The $\widehat{\mathbf{u}}_{k+1}$ is calculated based on a single step of the simulated dynamics from the ground truth solution $\mathbf{u}_{k}$. Assume $\mathbf{u}_{k}\in V$, one has
$\left\|\widehat{\mathbf{u}}_{k+1}-P_{V}\widehat{\mathbf{u}}_{k+1}\right\|_{2}=\Delta t\left\|P_{V^{\perp}}(A\mathbf{u}_{k}+B\widehat{C}\mathbf{u}_{k})\right\|_{2},$ (21)
and we may penalize the right hand side to promote trajectory alignment with the data manifold. Wrapping up all the ingredients, we state the loss function in the linear case as
$\min_{C}l(C):=\min_{C}\mathbb{E}_{(\mathbf{u},\mathbf{y})}\left(\left\|\mathbf{y}-C\mathbf{u}\right\|_{2}^{2}+\lambda\left\|P_{V^{\perp}}(A+BC)\mathbf{u}\right\|_{2}^{2}\right),$ (22)
and here $\lambda$ is a penalty strength parameter which can be chosen during training. We briefly discuss in the supplementary materials LABEL:lambda how to choose $\lambda$.
In above discussion, we have assumed that the underlying data subspace $V$ is known a priori or we can calculate $P_{V^{\perp}}$ directly. However, in most cases the only information we have is a set of resolved state variables $\mathbf{u}_{i}$s. Then, we can use standard dimension reduction methods such as principle component analysis [20] to obtain a subspace $V$, which can be thought as an approximation of the data manifold. The loss objective Eq. 22 can thus be calculated.
3.2 Algorithm in the general form
The discussion above applies to linear dynamics where the underlying data manifold is also linear. In this subsection we generalize it to non-linear cases and state the general form of our algorithm.
Let us return to our original formulation of the task Eq. 1, where we have a dataset $\left\{(\mathbf{u}_{1},\mathbf{y}_{1},t_{1}),(\mathbf{u}_{2},\mathbf{y}_{2},t_{2}),\cdots,(\mathbf{u}_{N},\mathbf{y}_{N},t_{N})\right\}$. As before, we assume that the resolved variable lies on a low-dimensional manifold $\mathcal{M}$ in the configuration space and the estimator $\phi_{\theta}(\mathbf{u},t)$ achieves high accuracy along this whole manifold while we have no guarantee outside the manifold. The key difference between the general case and the linear case is that this manifold may not be linear and we do not have an orthogonal decomposition.
To resolve this, we leverage an autoencoder to represent the data manifold structure, i.e. we learn two neural networks $E:\mathbb{R}^{m}\rightarrow\mathbb{R}^{l}$ (encoder) and $D:\mathbb{R}^{l}\rightarrow\mathbb{R}^{m}$ (decoder) to represent the manifold. The encoder maps the state variable $\mathbf{u}$ into a low dimensional space and decoder maps them back to original space. We optimize $E,D$ simultaneously to minimize the reconstruction error
$\min_{E,D}\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{u}_{i}-D(E(\mathbf{u}_{i}))\right\|_{2}^{2}.$ (23)
Here, we only use the resolved states $\mathbf{u}_{i}$ from the dataset. After training the autoencoder, we obtain a function
$F:\mathbb{R}^{m}\rightarrow\mathbb{R}_{\geq 0},\quad F(\mathbf{u})=\left\|\mathbf{u}-D(E(\mathbf{u}))\right\|_{2}^{2},$ (24)
which implicitly parametrizes the data manifold via $F(\mathbf{u})=0,\forall\mathbf{u}\in\mathcal{M}$. Notice that the linear version of autoencoder is exactly the principle component analysis we mentioned to identify linear subspace $V$. Such techniques to identify data manifolds have been used in improving robustness against adversarial attacks [5, 6]. Besides parametrizing the data manifold, $F$ can be used as an indicator of distribution shift, i.e. for any state variable $\mathbf{u}$, we use $F(\mathbf{u})=\left\|\mathbf{u}-D(E(\mathbf{u}))\right\|_{2}^{2}$ to measure its distribution shift w.r.t. the training data. This is used in all the figures in this paper to quantify distribution shift. Strictly speaking, training only ensures that the learned manifold is contained in the zero level-set of $F$. Thus, the gradient of $F$ belongs to the normal space of the manifold $\mathcal{M}$ at the point $\mathbf{u}$, i.e. $\nabla F\in\mathcal{M}^{\perp}$. Minimizing the inner product with $\nabla F$ now means partially minimizing the orthogonal components, which is the cause of the distribution shift in our formulation. This strictly generalizes the linear case Eq. 22, where the data subspace $V$ is given by $\operatorname{span}\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{l}\}$. Augmenting this basis to obtain an orthonormal basis $\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{l},\mathbf{v}_{l+1},\cdots,\mathbf{v}_{m}\}$ for the whole space, a choice of the function $F$ is given by
$F(\mathbf{u})=\frac{1}{2}\sum_{i=l+1}^{m}(\mathbf{u}^{T}\mathbf{v}_{i})^{2},$ (25)
and the gradient of $F$ is
$\nabla F(\mathbf{u})=\left(\sum_{i=l+1}^{m}\mathbf{v}_{i}\mathbf{v}_{i}^{T}\right)\mathbf{u},$ (26)
which is just the orthogonal projection onto $V^{\perp}=\operatorname{span}\{\mathbf{v}_{l+1},\mathbf{v}_{l+2},\cdots,\mathbf{v}_{m}\}.$ We can state the regularized loss function of our algorithm in general case:
$\min_{\theta}l(\theta):=\min_{\phi_{\theta}}\frac{1}{N}\sum_{i=1}^{N}\left(\left\|\mathbf{y}_{i}-\phi_{\theta}(\mathbf{u}_{i})\right\|_{2}^{2}+\lambda\left(\left(\nabla F(\mathbf{u}_{i})\right)^{T}\mathcal{L}(\mathbf{u}_{i},\phi_{\theta}(\mathbf{u}_{i}),t_{i})\right)^{2}\right).$ (27)
The first term above is the least squares term and the second term measures how far the estimated tangent vector $\partial_{t}\mathbf{u}_{i}=\mathcal{L}(\mathbf{u}_{i},\phi_{\theta}(\mathbf{u}_{i}),t_{i})$ deviates from the tangent space of $\mathcal{M}$ at $\mathbf{u}_{i}$. In the second term, $F(\mathbf{u})=\left\|\mathbf{u}-D(E(\mathbf{u}))\right\|_{2}^{2}$ and $D,E$ are optimized via Eq. 23 and frozen during the training of $\phi_{\theta}$. Since the loss function is regularized by a term which approximately measures the deviation of the tangent vector $\mathcal{L}(\mathbf{u}_{i},\phi_{\theta}(\mathbf{u}_{i}),t_{i})$ from the tangent space of the data manifold, we name our algorithm “MLHS with tangent-space regularized estimator”. The overall algorithm can be summarized in Algorithm 1.
Algorithm 1 MLHS with tangent-space regularized estimator
0:  $\left\{(\mathbf{u}_{1},\mathbf{y}_{1},t_{1}),(\mathbf{u}_{2},\mathbf{y}_{2},t_{2}),\cdots,(\mathbf{u}_{N},\mathbf{y}_{N},t_{N})\right\}$, resolved model, penalty strength $\lambda$.
1:  Learn a parameterized model which encode the structure of training data, i.e. $F_{\eta}(\mathbf{u})\geq 0,F_{\eta}(\mathbf{u}_{k})=0.$
2:  Freeze the parameters of this learned model.
3:  Introduce another surrogate model $\phi_{\theta}(\mathbf{u})$.
4:  for $k=1,2,\cdots,N$ do
5:     Predict the control variable $\widehat{\mathbf{y}}_{k}=\phi_{\theta}(\mathbf{u}_{k})$.
6:     Calculate the state variable after one-step iteration, i.e. $\widehat{\mathbf{u}}_{k+1}=\mathbf{u}_{k}+\Delta t\mathcal{L}(\mathbf{u}_{k},\widehat{\mathbf{y}}_{k},t_{k})$.
7:     Form the loss $l(\theta)=\mathbb{E}\left[\left\|\mathbf{y}_{k}-\widehat{\mathbf{y}}_{k}\right\|_{2}^{2}+\lambda\left(\left(\nabla F(\mathbf{u}_{k})\right)^{T}\mathcal{L}(\mathbf{u}_{k},\widehat{\mathbf{y}}_{k}),t_{k})\right)^{2}\right]$.
8:     Backpropogate to update $\theta$.
9:  end for
9:  tangent-space regularized estimator: $\phi_{\theta}$.
Comparing to literature of distributionally robust optimization (DRO) [40] and numerical analysis of dynamical systems [48, 47], our work takes advantages from both. First, the tangent-space regularized estimator belongs to the family of DRO in the sense that it minimizes the data-driven module of the unresolved part over some perturbations of the data distribution. However, unlike the original DRO which considers all possible perturbations of the data distribution in its neighborhood under some metric, e.g. Wasserstein metric  [3], our regularization focus on those perturbations caused by the resolved components of the dynamics. These facilitate us to take advantage of the partial knowledge of the dynamics. As a result, the regularizer we proposed is more problem-specific Moreover, from the perspective of DRO, our method takes into consideration the behavior of estimator under certain perturbation normal to the data manifold, which is much more tractable than guarding against arbitrary perturbations in distribution space. In Section 4, we will implement other benchmarks with general regularizations and compare with our algorithm to illustrate the benefit of combining resolved model information into regularization.
Furthermore, we can understand our method through the lens of the stability in numerical analysis of differential equations. The classical result that the least squares estimator converges to the ground truth in the limit of large datasets can be understood as a consistency statement. Here, we show that this is insufficient, and in order to ensure convergence of the simulated trajectories one needs to also promote stability, and our proposed regularizer serves precisely this role. This regularization approach should be contrasted with recent works [54, 7, 25, 16] on learning dynamical models from data which build stability by specifying the model architecture.
3.3 Interpretation and analysis of the algorithm
Let us now quantify in the linear case, the gains of our algorithm over the ordinary least squares estimator. We first calculates the exact formula for two estimators. Due to the low rank structure, OLS algorithm allows infinitely many solutions. Throughout this section, we make the common choice of the minimum 2-norm solution.
Proposition 3.1.
Consider the linear dynamics Eq. 12, the training data of the variable $\mathbf{u}$ is arranged into a data matrix $\mathbf{U}\in\mathbb{R}^{m\times N}$. Moreover, we assume the column space of $\mathbf{U}$ is contained in a subspace $V\subset\mathbb{R}^{m}$ with associated projection operator $P_{V}$. Then, the OLS estimator and tangent-space regularized estimator for the unresolved component $Y=C^{*}\mathbf{U}+\epsilon$ have the following form:
$\widehat{C}_{OLS}=$ $\ C^{*}P_{V}+\epsilon\mathbf{U}^{\dagger},$ (28)
$\widehat{C}_{TR}=$ $\ (\mathbf{I}+\lambda B^{T}P_{V^{\perp}}B)^{-1}(C^{*}P_{V}+\epsilon\mathbf{U}^{\dagger}-\lambda B^{T}P_{V^{\perp}}AP_{V}).$
Specifically, in the noiseless scenarios, $\widehat{C}_{OLS}$ recovers $C^{*}$ on the subspace $V$ but vanishes on its orthogonal complement. In order to interpret our estimator, let us consider a simpler case with $B=\mathbf{I},A=\mathbf{0}$. This corresponds to the dynamics
$\partial_{t}\mathbf{u}=C^{*}\mathbf{u},$ (29)
and the regularized estimator is given by
$\widehat{C}_{TR}=(\mathbf{I}+\lambda P_{V^{\perp}})^{-1}(C^{*}P_{V}+\epsilon\mathbf{U}^{\dagger})=(\mathbf{I}+\lambda P_{V^{\perp}})^{-1}\widehat{C}_{OLS}.$ (30)
Therefore, our estimator performs a weighted least squares regression which penalizes the direction perpendicular to the data subspace $V$, as controlled by $\lambda$. As the error along this normal direction is the main cause of the distribution shift, it is sensible to reduce it more than the error along the tangent direction, which cause deviations inside the manifold. This shows that the tangent-space regularized estimator does in fact mitigate distribution shift.
Next, we move on to analyze the accuracy gain of the whole simulation algorithm by proving a bound for the error of the simulation trajectory. Denote the statistical error of tangent-space regularized estimator by
$l_{TR}(\widehat{C})=\mathbb{E}\left(\left\|(C^{*}-\widehat{C})\mathbf{u}\right\|_{2}^{2}+\lambda\left\|P_{V^{\perp}}(A+B\widehat{C})\mathbf{u}\right\|_{2}^{2}\right)$ (31)
We define the following function which helps simplify the notation
$Q(r,T)=\frac{e^{rT}-1}{r},\quad r,T>0.$ (32)
Theorem 3.2.
Under the same setting as Proposition 3.1, assume both the OLS estimator and TR estimator $\widehat{C}_{OLS},\widehat{C}_{TR}$ are optimized to have error bounded by $\delta$ so that $l_{OLS}(\widehat{C}_{OLS}),l_{TR}(\widehat{C}_{TR})<\delta$ in Eq. 15 and Eq. 31. Suppose the true trajectory is simulated from an initial condition $\mathbf{u}(0)\in V$ which follows the same distribution of the training data. Then, the errors of OLS and our algorithm are bounded respectively by
$\mathbb{E}\left\|\widehat{\mathbf{u}}_{TR}(T)-\mathbf{u}(T)\right\|^{2}\leq$ $\delta Q(\left\|A+B\widehat{C}_{TR}P_{V}\right\|^{2}+3,T)$ (33)
×$\times\left(\left\|B\right\|^{2}+\frac{Q(\left\|P_{V^{\perp}}(A+B\widehat{C}_{TR})\right\|^{2}+2,T)}{\lambda}\right),$
$\mathbb{E}\left\|\widehat{\mathbf{u}}_{OLS}(T)-\mathbf{u}(T)\right\|^{2}$ $\leq\delta Q(\left\|A+B\widehat{C}_{OLS}P_{V}\right\|^{2}+\left\|B\widehat{C}_{OLS}\right\|^{2}+3,T)\left\|B\right\|^{2}.$
The proof idea is illustrated in deriving Eq. 19 and we leave step-by-step derivations in the supplementary materials Section 3. To understand this result, let us focus on the exponential part of two error bounds as this will dominate the error when the simulation time $t$ is large. The exponent of our algorithm is bounded by $(\left\|A+B\widehat{C}_{TR}P_{V}\right\|^{2}+3)t$ provided that the penalty parameter $\lambda$ is large enough. This is related to the intrinsic stability constant associated with the dynamics, underlying manifold. If we assume that two terms $\left\|A+B\widehat{C}_{OLS}P_{V}\right\|,\left\|A+B\widehat{C}_{TR}P_{V}\right\|$ are close to each other, this suggests that our algorithm may have a slower error accumulation rate for large $t$. As the training data lies on a subspace $V$ of the whole state space, we do not have any bound on the behavior of $\widehat{C}_{OLS}$ outside this subspace. For example, the orthogonal complement $V^{\perp}$ may contain an eigenspace of $B\widehat{C}_{OLS}$ with a large eigenvalue. If this is the case, $\left\|B\widehat{C}_{OLS}\right\|$ may be significantly larger than $\left\|A+B\widehat{C}_{TR}P_{V}\right\|$, and the improvement of our algorithm may be more promising. In supplementary material LABEL:fig:linear-cmp, we provide a toy example which illustrates the effectiveness of our method comparing to OLS.
4 Numerical experiments
In this section, we present numerical experiments which illustrate the effect of the tangent-space regularized algorithm presented in Algorithm 1. We use the experiments to illustrate two phenomena. The first demonstrates the effectiveness of our algorithm, for which we test on the reaction-diffusion equation and the incompressible Navier-Stokes equation, and compare with benchmark algorithms using least squares, as well as simple regularizations that do not account for the structure of the resolved dynamics or the data manifold. For the second goal, we show that our method brings more significant improvements in problems with severe distribution shifts. This is demonstrated on the same family of PDEs, i.e. reaction-diffusion equation with varying diffusion coefficient and Navier-Stokes equation with different Reynolds numbers. These varying parameters can be understood - as we show numerically - as quantitative indicators of the severity of distribution shift. The implementation of our method and experiment reproduction is found in the repository [57].
4.1 Tangent-space regularized algorithm improves simulation accuracy
We first show that the proposed algorithm can be advantageous on several prototypical machine learning augmented scientific computing applications.
In the first experiment, we use the FitzHugh-Nagumo reaction-diffusion Eq. 11 equation as the test case of our algorithm. The resolved-unresolved structure is the coarse-fine grid correction structure discussed in Section 2.2. This helps us obtain high-fidelity simulation results by combining a low-fidelity numerical solver with a fast neural network surrogate model for correction. Here we choose $\alpha=0.01,\beta=1.0,D=\begin{pmatrix}\gamma&0\\ 0&2\gamma\end{pmatrix},\gamma\in\{0.05,0.10,0.15,0.20,0.25\}$ in the simulation which guarantee the existence of Turing patterns. The initial condition $\mathbf{u}_{0}$ is generated by i.i.d. sampling from a normal distribution:
$u(x,y,0),v(x,y,0)\sim\mathcal{N}(0,1).$ (34)
This system is known to have a dynamical spatial patterns where initial smaller “clusters” gradually transforms and merges to form bigger cluster according to the balance between diffusion and reaction effects [33]. In literature, the importance of the sampling procedure is also emphasized and various adaptive sampling method based on error and particle methods are studied in [12, 52, 58].
Now, we explain how the training data is prepared. Recall that in Eq. 6 we discussed the hybrid simulation structure for a correction procedure under a pair of coarse and fine mesh:
[…]
where the mapping $\phi$ between the fine grid state $\mathbf{u}_{k}^{2n}$ to the correction term $\mathbf{y}_{k}^{2n}$ is the unresolved model we desire to learn. Here $I_{n}^{2n},R_{2n}^{n}$ are fixed restriction and interpolation operators between two meshes, i.e.
$R_{2n}^{n}:\mathbb{R}^{2n\times 2n}\rightarrow\mathbb{R}^{n\times n},\quad\mathbf{u}_{ij}^{n}=\frac{\mathbf{u}_{2i,2j}^{2n}+\mathbf{u}_{2i,2j+1}^{2n}+\mathbf{u}_{2i+1,2j}^{2n}+\mathbf{u}_{2i+1,2j+1}^{2n}}{4},$ (36)
$I_{2n}^{n}:\mathbb{R}^{n\times n}\rightarrow\mathbb{R}^{2n\times 2n},\quad\mathbf{u}_{2i,2j}^{2n}=\mathbf{u}_{2i,2j+1}^{2n}=\mathbf{u}_{2i+1,2j}^{2n}=\mathbf{u}_{2i+1,2j+1}^{2n}=\mathbf{u}_{ij}^{n}.$
To obtain training data for $\mathbf{u}_{n}^{t},\mathbf{u}_{2n}^{t}$, we simulate Eq. 11 (with the same initial condition) on two meshes of sizes $64\times 64$ and $128\times 128$ respectively, with time step $\Delta t=0.01,N_{\text{step}}=T/\Delta t=100$, using a second order Crank-Nicolson scheme. It remains to calculate the labels for the training data, i.e. $\mathbf{y}_{k}^{2n}=\phi(\mathbf{u}_{k}^{2n})$. These are obtained by using the first half of the Eq. 36, i.e.
$\mathbf{y}_{n}^{t}=\mathbf{u}_{2n}^{t+1}-I_{n}^{2n}\circ f_{n}(R_{2n}^{n}(\mathbf{u}_{2n}^{t}))),$ (37)
Finally, we pair them together to form the training data $\left\{(\mathbf{u}_{n}^{1},\mathbf{y}_{n}^{1}),\cdots,(\mathbf{u}_{n}^{N},\mathbf{y}_{n}^{N})\right\}$. The data here is different from the one introduced in Section 2.3 since the system is autonomous, so we do not include time indices. To summarize, the training data contains two patches of field component patterns of size $100\times 128\times 128$ corresponding to $\mathbf{u},\mathbf{v}$ respectively and two patches $\mathbf{y}_{u},\mathbf{y}_{v}$ of the same shape and generated according to Eq. 36.
The baseline algorithms we implemented to compare with ours is $\phi_{\text{OLS}}$, which minimizes the mean-squared error on the training data, i.e.
$\phi_{\text{OLS}}=\arg\min\mathbb{E}_{\mathbf{u}\sim\rho_{0}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|_{2}^{2},$ (38)
where we use $\rho_{0}$ to denote the distribution of the field variable generated by certain simulation condition Eq. 34. The corresponding numerical results are labelled by “OLS”.
Figure 3: Simulation of the reaction-diffusion equation and configurations at time step: 0, 100, 200, 300, 500, (a) ground truth; (b) simulated fluid field using OLS estimator; (c) simulated fluid field using TR estimator. (d) is the comparison of the ordinary least squares and regularized one on reaction-diffusion equation: the solid line represents the averaged error along the trajectory and the dash line represented the averaged distribution shift calculated using Eq. 24. The shadow part is the standard deviation in ten random experiments with different training and testing data. Both the error and the distribution shift increase at a lower speed in TR simulation than in OLS simulation. This suggests that the better performance of TR algorithm in reaction-diffusion equation is partially caused by improving the distribution shift issue.
We summarize the comparison results in Fig. 3. Here, solid lines represent the errors along the trajectories and dashed lines represent the distribution shift. The shaded part is the standard deviation over ten random experiments with different split of the training and testing data. As can be observed, compared to our method the OLS baseline incur greater errors and distribution shifts. Moreover, this difference increases with time. In contrast, consistent with our analysis in the simplified setting, our algorithm mitigates both the distribution shift and the trajectory error.
Next, we consider the Navier-Stokes equation Eq. 3 as the test case of our algorithm Algorithm 1. In this case, the resolved model is an explicit scheme to evolve the velocity field while unresolved model is calculated via a fast CNN surrogate, which can accelerate the simulation for large grid system. The simulation domain is rectangular with aspect ratio $1/4$. On the upper and lower boundary we impose no-slip boundary condition on the velocity and on the inlet, i.e. $x=0$ we specify the velocity field $\mathbf{u}$. Moreover, at the outlet, i.e. $x=n$ we pose zero-gradient conditions on both the horizontal velocity and pressure, i.e. $\frac{\partial p}{\partial n}=\frac{\partial u}{\partial n}=0$. The inflow boundary is defined according to following functional forms:
$\mathbf{u}(0,y,t)=\begin{pmatrix}u(0,y,t)\\ v(0,y,y)\end{pmatrix}=\begin{pmatrix}\exp\{-50(y-y_{0})^{2}\}\\ \sin t\cdot\exp\{-50(y-y_{0})^{2}\}\end{pmatrix}$ (39)
The training configuration is calculated using the projection method [10] with staggered grid, i.e. the pressure is placed at the center of each cell and horizontal (vertical) velocity is placed at vertical (horizontal) edge of each cell. The jet location $y_{0}$ is selected from $0.3$ to $0.7$ uniformly.
Again, we compare the performance of our algorithm and the OLS method. As illustrated in Fig. 4, naively using OLS to estimate the unresolved models will generally lead to error blow-up, whereas using our regularized estimator we can obtain a reasonable simulation along the whole time period. Moreover, the trend of the trajectory error fits that of the distribution shift. In particular, the error and distribution shift blow-ups occur simultaneously. Our algorithm again decreases the trajectory error by mitigating the problem of distribution shift.
Figure 4: Simulation of the Navier-Stokes equation and configurations at time step: 0, 60, 120, 180, 240, (a) ground truth; (b) simulated fluid velocity field using OLS estimator; (c) simulated fluid velocity field using TR estimator; (d) the comparison of the OLS estimator and the regularized one. The solid lines represent the trajectory error and the dashed lines represent distribution shift calculated using Eq. 24. We observe that the OLS estimator leads to error (and DS) blow-up at around 270 time steps, while our method remains stable. Again, the trends of the error and the distribution shift are highly correlated, and our method that controls distribution shift leads to error control, consistent with our analyses.
The machine-learning augmented NS simulations appear to have a greater degree of distribution shift, and our method has a relatively stronger benefit in this case. This suggests that our method is most beneficial in problems where the intrinsic distribution shift is severe. In the next subsection, we will study this more systematically by testing our method on the reaction-diffusion equations with varying diffusion coefficient $\gamma$ and Navier-Stokes equations with varying Reynolds numbers.
4.2 Performance under distribution shift of different magnitudes
We now demonstrate concretely that our method brings more significant improvements when the intrinsic distribution shift is more severe. Intuitively speaking, since our algorithm penalizes the deviation of the simulated trajectories to the data manifold while OLS totally ignores this, it would be reasonable to expect that our algorithm will outperform OLS more in the problems where such deviations are large. A sanity check showing different distribution shift extents under different parameters is provided in the supplementary materials LABEL:fig:NSRe-ds.
Here, we introduce two more baselines for ablation studies. Comparing with the OLS algorithm, these algorithms both add some regularization to the loss function while their regularizations are general and not dynamics-specific compared to our regularization. Consequently, any improved performance in our algorithm over these regularized algorithms suggests that targeted regularization, which takes into account the structure of the resolved part, effectively balances accuracy and stability.
The first baseline regularizes the original least squares objective by a term which quantifies some complexity the network model, e.g.
$\phi_{\text{mOLS}}=\arg\min\mathbb{E}_{\mathbf{u}\sim\rho_{0}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|_{2}^{2}+\lambda\left\|\theta\right\|_{2}^{2}.$ (40)
Here the naive $L^{2}$ regularized algorithm is implemented by adjusting the weight decay [31] parameter of the Adam solver during the training process. The corresponding numerical results are labelled by “mOLS” (“m” for modified).
The second baseline adds noise to the training data during training, which can be formulated as
$\phi_{\text{aOLS}}=\arg\min\mathbb{E}_{\mathbf{u}\sim\rho_{\xi}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|_{2}^{2}.$ (41)
Notice that the training data follows the distribution $\rho_{\epsilon}=\rho_{0}*\xi$ where *** denotes the convolution operator and $\xi$ is some random noise which we take to be Gaussian random variable with an adjustable variance. In implementation, we generate the noise and add to the input during each epoch so the inputs are not the same across epochs. Moreover, this noise $\xi$ should be distinguished from the observational noise $\epsilon$ we introduced in Section 3. This method is used in several previous work [46, 36] to improve stability and prevent the error from accumulating too fast. As this technique is similar to adversarial attack and data augmentation [44, 14, 13], we name this estimator and algorithm as “aOLS” (“a” stands for adversarial). We slightly abuse the terminology here by naming this random perturbation as some kind of adversarial attack.
Table 1: Performance on reaction-diffusion equation over 10 experiments.
$\gamma=0.05$ $\gamma=0.10$ $\gamma=0.15$ $\gamma=0.20$ $\gamma=0.25$
grid size: $64\times 64$
OLS 8.40e+01 1.04e+01 8.99e+00 2.22e+00 6.78e-01
mOLS 8.90e+00 9.22e+00 1.41e+00 2.26e+00 9.98e-01
aOLS 3.57e+01 4.00e+01 2.40e+00 1.79e+00 1.06e+00
TR 1.23e-01 1.06e-01 1.38e-01 1.85e-01 2.15e-01
Diff 99.9% 99.0% 98.5% 91.7% 68.3%
grid size: $128\times 128$
OLS 4.10e+02 8.18e+01 5.11e+01 4.88e+00 2.25e+00
mOLS 2.40e+01 3.20e+01 2.20e+00 1.07e+01 1.25e+00
aOLS 1.46e+02 2.58e+02 8.69e+00 3.30e+00 1.79e+00
TR 2.79e-01 1.66e-01 2.42e-01 1.81e-01 1.91e-01
Diff 99.9% 99.8% 99.5% 96.3% 91.5%
In Table 1, we test the algorithms under different diffusion coefficient $\gamma=0.05$, $0.10,0.15,0.20,0.25$. We calculate the relative trajectory error at time $T=1000$, i.e. $\frac{\left\|\mathbf{u}_{k}-\widehat{\mathbf{u}}_{k}\right\|}{\left\|\mathbf{u}_{k}\right\|}$ over ten random split of the training and test data. The last row calculates the performance difference between our method and the OLS baseline. We observe that under almost all the settings our method out-performs OLS, mOLS, aOLS baselines, and comparing horizontally, we see that the relative improvement decreases with $\gamma$, as diffusion term mitigate the distribution shift.
We now consider the Navier-Stokes equation. Since the baseline method quickly leads to error blow up, we use another comparison criterion. We define a stopping time $t_{K}=\arg\max_{t}\frac{\left\|\mathbf{u}_{t}-\widehat{\mathbf{u}}_{t}\right\|}{\left\|\mathbf{u}_{t}\right\|}\leq K$ where $K$ is a error threshold to be determined during the experiments. We calculate the first time the trajectory error reaches a threshold under different Reynolds numbers and different mesh sizes. The results are shown in Table 2. Under almost all the scenarios, our method outperforms all the baseline algorithms by a larger $t_{K}$. This validates the effectiveness of our algorithm. Comparing Table 2 horizontally, one finds that the improvement of our method is also increasing with respect to the Reynolds number. As the flow field generally becomes more complex (and sensitive) with increasing Reynolds number, our method is expected to bring bigger improvements. This is consistent with our experiments.
Table 2: Comparison of stopping time $t_{K}$ on Navier-Stokes equation over 10 experiments.
$Re=100$ $Re=200$ $Re=300$ $Re=400$ $Re=500$
grid size: $128\times 32$
OLS $435\pm 15$ $301\pm 20$ $240\pm 12$ $177\pm 17$ $120\pm 7$
mOLS $420\pm 12$ $333\pm 14$ $254\pm 12$ $204\pm 9$ $210\pm 7$
aOLS $477\pm 25$ $392\pm 31$ $450\pm 23$ $239\pm 25$ $321\pm 20$
TR $\mathbf{506\pm 18}$ $\mathbf{497\pm 28}$ $\mathbf{482\pm 29}$ $\mathbf{442\pm 21}$ $\mathbf{452\pm 13}$
grid size: $256\times 64$
OLS $602\pm 38$ $571\pm 35$ $355\pm 25$ $301\pm 11$ $148\pm 22$
mOLS $623\pm 18$ $531\pm 26$ $489\pm 20$ $417\pm 10$ $298\pm 9$
aOLS $\mathbf{651\pm 48}$ $601\pm 42$ $\mathbf{595\pm 39}$ $378\pm 31$ $398\pm 32$
TR $636\pm 35$ $\mathbf{602\pm 28}$ $585\pm 23$ $\mathbf{422\pm 35}$ $\mathbf{402\pm 28}$
5 Conclusion
In this paper, we establish a theoretical framework for machine-learning augmented hybrid simulation problems, where a data-driven surrogate is used to accelerate traditional simulation methods. We identify the cause and effect of distribution shift, i.e. the empirically observed phenomenon that the simulated dynamics may be driven away from the support of the training data due to systematic errors introduced by the data-driven surrogate, magnified by the resolved components of the dynamics. Based on this, we propose a tangent-space regularized algorithm for training the surrogate for the unresolved part, which incorporates the resolved model information to control deviations from the true data manifold. In the case of linear dynamics, we show our algorithm is provably better than the traditional training method based on ordinary least squares. Then, we validate our algorithm in numerical experiments, including Turing instabilities in reaction-diffusion and fluid flow. In both cases, our method outperforms baselines by better mitigating distribution shift, thus reducing trajectory error.
Acknowledgements
The research work presented is supported by the National Research Foundation, Singapore, under the NRF fellowship (project No. NRF-NRFF13-2021-0005).